You are the financial planner for Johnson Controls. Assume last year’s profits were $710,000. The board of directors decided to forgo dividends to stockholders and retire high-interest outstanding bonds that were issued 5 years ago at a face value of $1,330,000. You have been asked to invest the profits in a bank. The board must know how much money you will need from the profits earned to retire the bonds in 10 years. Bank A pays 6% compounded quarterly, and Bank B pays 7% compounded annually.(use Table 1 and Table 2 provided.) (Do not round intermediate calculations. Round your answers to the nearest dollar amount.)

a-1. Which bank would you recommend?

Bank A
Bank B

a-2. How much of the company’s profit should be placed in the bank?

Profit            $

b. If you recommended that the remaining money not be distributed to stockholders but be placed in Bank B, how much would the remaining money be worth in 10 years?

Future Value            $
Answer #1

Effective Annual Return = (1+Periodic Return)^Number of periods – 1

Bank A = (1+6%/4)^4 – 1

= 6.1364%

Bank B = 7%

Hence, Bank B is recommended, since it is offering higher rate of interest

a-2 Profit required to be placed = Present value of amount required in future

= 1,330,000/(1.07)10

= $676,104.56

b.Remaining money = 710,000-676,104.56 = $33,895.44

Amount in 10 years = 33,895.44(1.07)10 = $66,677.46